This page has already been transcribed. You can find new pages to transcribe here.

Transcription

[Page 158]

leave for the Queensland Boys before Christmas.  Another of my work mates in the Stokehole in the Emden fight was made Leading Stoker (W. Wyatt) to day, good Luck to him for he has a hard fight for it. I am now the Lone Star of Chili who has not got made nor am I likely not being interested in it.  (No Press News.)  We are trying to work up another Magazine.  An issue was made in Slop today (Straw Hats).  We are going to try a new experiment wood and Oil fuel.  A Water Snake was caught in the first cutter this morning about 1 ft. 6in. long and put in a Bottle.  We are now pested by swarms of flys from the Swamps.  The Purser has got no soap in the Stores nor tobacco so things are looking bad.  One half of the Ships Company is suffering with Prickly Heat and the other half with Boils and we are all fed up with being amongst the Black race, and in the Tropics.

12th.
Division, Church.  61 Men under Mr. Dennis went wood chopping again. The Divers went down to clean the under water fittings and clear the water line.  There is a concert to night at 7.45 PM, Weather permitting.  There was a disturbance in the Band owing to the Election of a Drum Major but that has been washed out until a later date.  Make and Mend Clothes from 1.30 PM.  L. Duty Pulse 96.  Took a slide of Blood from my finger tip, Dr. Marne.  Leave was granted from 1.30 PM to 6.30 PM to the Port Watch for Bathing.  12 Months Agoney completed, how much more.  Mr. Hodson went fishing in a Schooner.  Wet weather concert prosponed until tomorrow, Weather Permitting.

13th.
​At 7 AM Mr. Hodson was brought on board with his left fore arm nearly blown off and his chest was all burnt with a dymanite explosion owing to a defective detonator which he was going to use to blow up the fish.  A.B. Giles LTO was hurt in the Stomach at the same time.  The Doctors operated on Mr. Hodson at nine AM and are still working 1 PM.  Did not see the doctor to day owing to the operation.

Current Status: 
Completed